NPP Projection Report 1/24 Open


To do this evening/ was completed: Gil radioed me about no image at the start of the first round in T2. I checked the VNC viewer for the touch panel and the lamp was off, no "unexpected light off" warning so it was never struck (even though there was a lamp on cue in the playlist). Logs confirmed the cue just didn't go through--there was no attempt to strike the lamp. Please check to make sure the lamp shuts off at the end of the night.

Presentation Issues:

Preshow Clips/Playlist: I was also radioed about trailers in T5 being too quiet for the first round (this was the earliest first run show of the day). I checked the playlist and there was a cue to lower the volume for the book promo (as there should be), but no cue to raise it back for the first run trailers. This was the case for most playlists for this week and next, though some were correct. I updated cues on all of those which needed it.

Converted and ingested Desmond's slides for Feb (Snow Day, Unbreakable, Die Hard With A Vengeance)

Trailer Updates:

Special Events/ AV set up: The Birds went nicely in T1. Print is rewound for tomorrow.

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content): Downloaded all the NoBudge shorts in a folder on the desktop ("NoBudge 49"). Started DCP-ing, first one is done and the second is currently converting. Going into that same NoBudge folder on the desktop.

Outgoing/Shipping (content):

Print/video Inspection:

Asset tracking doc updated/checked?:

Maintenance/cleaning: Added some oil to the P1 motor in T1. It was getting low.

Cue sheets updated/checked (every shift):

Marquee updated?:

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will

NPP Projection Admin 7/6